OpenStudy (anonymous):

What is the solution of the equation 9 = -2x + 15?

OpenStudy (anonymous):

Alright you want to isolate the coefficient and x one side and get all constants (numbers without variables x,y,z,etc. Using The subtraction equality theorem, that states that if you subtract on one side , you must subtract the same amount on the other side to keep the equation balanced if you want the 15 gone from the right side, you can subtract by 15, but you have to subtract 15 to the left also so you get something like this \[9-15=-2x+15-15\] \[9-15=-2x+0\]

OpenStudy (anonymous):

9 = -2x + 15 -15 -15 -6 = -2x -- -- 2 2 -3=x

OpenStudy (anonymous):

9-15 is simply -6 \[-6=-2x\] now you need to use the division equality theorem which states that if you divide one side by a number, you have to do it also to the other side

OpenStudy (anonymous):

i think my work is well put what do you think @SJOJ4552

OpenStudy (anonymous):

so if you want x by itself you have to divide by a -2 on the right side because \[\frac{-2}{-2}=1\] but since you divide by -2 on the right side you also have to do it to the left to balance it \[\frac{-6}{-2}=\frac{-2}{-2}x=x\]

OpenStudy (anonymous):

your answer is wrong

OpenStudy (anonymous):

really ??

OpenStudy (anonymous):

you lost a negative \[-3=-x\neq x\]

OpenStudy (anonymous):

yup so positive 3 opps lol but still done right